The silicon controlled rectifier connected device has 100 sin 314t supply voltage. The resistance and inductance are connected in series with the silicon controlled rectifier.If the latching current 4mA, find the minimum width of the gate trigger pulse required to properly turning ON the SCR?


1
Expert's answer
2021-11-21T17:23:25-0500

"V=V_0sin(wt)\\V=100sin(314t)"

"V_0=100V \\I=4mA\\L=0.2mH"

We know that

"E=L\frac{di}{dt}"

Now

"I=\frac{V_0}{L}t"

"t=\frac{LI}{V_0}"

Put value


"t=\frac{0.2\times4\times10^{-3}}{100}=8\times10^{-6}sec=8\mu sec"